Given utility function U=144q-1.5qsquared


Find marginal utility


Calculate the utility maximising quantity


What is marginal rate of substitution

1
Expert's answer
2021-11-16T13:37:45-0500

I) marginal utility;

"MU=\frac{\delta U}{\delta q}"


"=144-3q"


ii)Utility maximizing quantity

"q=\frac{144}{3}"


"=48"


iii) marginal rate of substitution

Marginal rate of substitution (MRS) is the amount of a good that a consumer is willing to consume compared to another good, as long as the new good is equally satisfying.